OpenStudy (anonymous):

the expression 10/(3 + i) is equivalent to: 1) 3 - i 2) (15 + 5i)/4 3) 3 + i 4) 5/4

OpenStudy (anonymous):

\[\frac{10}{3+i}=10\frac{1}{3+i}=10\frac{1}{3+i} \times \frac{3-i}{3-i}=10\frac{3-i}{3^2-i^2}=10\frac{3-i}{10}\]

OpenStudy (anonymous):

\[=3-i\]

OpenStudy (anonymous):

Your first available choice.